what are the solution of equation 2sinx + cosx = 3
a. only one solution 
​b. no solution 
​c. infinetiy many solution
​d. finitly many solution
 

Grade:12th pass

1 Answers

Soumendu Majumdar
5 years ago
Dear Student,
-1\leq sin x\leq 1   and    -1\leq cos x\leq 1
so
2sinx + cos x=3 is only possible if both sin x & cos x are equal to 1.
But sin x & cos x cannot be equal to 1 at the same time
so option (b) is correct.
Hope it helps!